# Rethinking Explainable Machine Learning as Applied Statistics

### v4

@@ Abstract
In the rapidly growing literature on explanation algorithms, it often remains unclear what precisely these algorithms are for and how they should be used. In this position paper, we argue for a novel and pragmatic perspective: Explainable machine learning needs to recognize its parallels with applied statistics. Concretely, explanations are statistics of high-dimensional functions, and we should think about them analogously to traditional statistical quantities. Among others, this implies that we must think carefully about the matter of interpretation, or how the explanations relate to intuitive questions that humans have about the world. The fact that this is scarcely being discussed in research papers is one of the main drawbacks of the current literature. Luckily, the analogy between explainable machine learning and applied statistics suggests fruitful ways for how research practices can be improved.
Explainable Machine Learning, Position Paper
@@ 1 Introduction
In this position paper, we argue for a novel perspective on explainable machine learning: The field needs to recognise its parallels with applied statistics. [^1] Applied statistics is the use of statistical theory, methods, and tools to analyze data and solve real-world problems (Wilcox & Rousselet, 2023; Angrist & Pischke, 2009; Cox & Donnelly, 2011; Cox, 2018; Franconeri et al., 2021). Indeed, explainable machine learning and applied statistics both aim to summarize the behaviour of high-dimensional objects — functions in explainability and datasets/probability distributions in statistics (Wasserman, 2004). Moreover, both fields study how relevant questions about the real world can be answered given such summaries.
@@
    The matter of it’s interpretation.
Mathematically, explanation algorithms have the form of a functional – a function of functions (Rudin, 1991). They take a function and map it to a simpler object, often a vector. One might say that the functional “explains” a function by reducing its complexity. However, not all dimension-reducing functionals have explanatory value. For this, the functional also needs to formalize an intuitive concept, a nuanced matter debated in the philosophy of science (Justus, 2012). If this is the case, then we say that the functional has an interpretation. Traditional statistical quantities like the $p$-value, confidence intervals and visualizations usually have an interpretation.
In the current literature on explainable machine learning, papers often introduce novel explanation algorithms without discussing their interpretation. This is problematic in itself because “explanations” without an interpretation miss the point. It is additionally problematic since historically, even relatively simple statistics that have an interpretation have frequently been misinterpreted, especially by applied scientists and decision-makers (Wasserstein & Lazar, 2016). Consequently, it is not surprising that the outputs of explanation algorithms have frequently been misinterpreted (Molnar et al., 2020).
@@
To be concise, this paper focuses on post-hoc explanation algorithms, which comprise the bulk of the research in explainable machine learning over the last decade. However, the analogy between explainable machine learning and applied statistics applies broadly. In the final Section 9, we briefly discuss how it applies to mechanistic interpretability.
@@ 2 Post-Hoc Explanation Algorithms are Statistics of Functions
Notation. $\mathcal{D}$ denotes a probability distribution on $\mathbb{R}^{d}$ and $f:\mathbb{R}^{d}\to\mathbb{R}$ a function. We let $\mathcal{P}(\mathbb{R}^{d})$ denote the space of all probability distributions on $\mathbb{R}^{d}$ and $\mathcal{F}(\mathbb{R}^{d})$ the space of all functions $f:\mathbb{R}^{d}\to\mathbb{R}$.
@@ 2.2 Explanation Algorithms: Statistics of functions
    SHAP values (Lundberg & Lee, 2017) $\mathnormal{F}_{i}(f,\mathcal{D},x)=$
    $\sum_{S\subseteq[d]\setminus\{i\}}{\frac{|S|!(d-|S|-1)!}{|d|!}}[\mathbb{E}_{\mathcal{D}}(f|x_{S\cup\{i\}})-\mathbb{E}_{\mathcal{D}}(f|x_{S})].$
@@
    The closest point to $x$ with a different label than $x$
    $\operatorname*{arg\,min}_{z\in\{y|f(y)!=f(x)\}}||x-z||_{1}.$
@@
In summary, Definition 2.3 captures most, if not all, local post-hoc explanation algorithms. It also captures various other properties of functions, such as fairness- and robustness metrics. Note that when we talk about an algorithm, we talk about the function implemented through that algorithm.
@@ 3 The Analogy Between Explainable Machine Learning and Applied Statistics
We have seen that post-hoc explanations and other properties of classifiers can be seen as statistics of functions. We have also seen that there is a formal similarity between statistics of functions and statistics of probability distributions or of datasets. This suggests that we should think about explanations and traditional statistical quantities along analogous lines and that explainable machine learning is statistics for high-dimensional functions.
Historically, researchers in statistics have studied how the properties of probability distributions and large datasets can be summarized using distribution and dataset statistics (Stigler, 2002). In explainable machine learning, researchers have studied how the properties of learned functions can be summarized using explanations. In both cases, the principle is the same: We take a high-dimensional object, map it into a low-dimensional space, and then use the resulting statistical value(s) to answer relevant questions that we are interested in.
As we argue in the following Section 4, the analogy between explainable machine learning and statistics extends toward the crucial question of how to interpret explanations in applications. Before this, we want to briefly discuss two insights from the mathematical framework.
@@
Insight 2: A large part of the literature in explainable machine learning studies the mathematical properties of statistics of functions.
The mathematical framework introduced in Section 2 can help us understand and categorize research questions in explainable machine learning. In particular, we can ask whether a research question can be fully expressed within the mathematical framework. In other words: What research in explainable machine learning studies the mathematical properties of high-dimensional functionals? It turns out that a large part of the literature studies such questions. For example, many papers study the sensitivity of a statistic $F$ with respect to $f$ and $x$ (Slack et al., 2020; Anders et al., 2020; Lakkaraju et al., 2020), or the behaviour of a statistic $F$ on a subset of functions $\mathcal{I}\subset\mathcal{F}(\mathbb{R}^{d})$ (Garreau & von Luxburg, 2020; Bordt & von Luxburg, 2023).
Research on the mathematical properties of statistics can be useful. In particular, research in statistics also has strong mathematical aspects (Larsen & Marx, 2005).
@@ 4 Statistics and their Interpretation
Both in applied statistics and explainable machine learning, the starting points are intuitive questions humans have about the real world. For example
@@ 4.1 Formalizing intuitive concepts
How can we say whether a given statistic has an interpretation? It turns out that the relationship between intuitive concepts and their formalizations has long been studied in philosophy, where this kind of problem is referred to as conceptual analysis (Chalmers, 1996; Chalmers & Jackson, 2001; Jackson, 1998; Justus, 2012). A central insight in this literature is that the intuitive concept is informal and pre-theoretic, whereas the formalized concept is theoretic and part of a workable theory (Carnap, 1945, 1950). Since the intuitive concept is not formalized, there is no strict way that we could prove that a statistic formalizes an intuitive concept. However, there are arguments for whether a formalization captures an intuitive concept that come close to proving such statements.
@@ 4.2 Statistics of functions that have an interpretation
Examples of post-hoc explanation methods that have an interpretation are counterfactual explanations (Wachter et al., 2017) and anchors (Ribeiro et al., 2018). Counterfactual explanations formalize the intuitive concept “What would need to change to get a different prediction?”. Anchors formalize the intuitive concept that “Similar observations should have the same label”. What is important is that for both counterfactual explanations and anchors, it is clear which intuitive questions can and cannot be answered given the explanations. For example, a counterfactual explanation applies only to an individual data point and does not necessarily tell us how the input features of other points would need to be changed. Similarly, anchors only apply to a specific region of the input space.
@@ 4.3 Many post-hoc explanations lack an interpretation
Take feature attribution methods as an example (Ribeiro et al., 2016; Lundberg & Lee, 2017; Selvaraju et al., 2017). These methods address the question, “Which features are important for the prediction?”. As such, there is an intuitive concept that the different methods attempt to formalize. However, the problem is that the relationship between formalism and intuitive concept is not sufficiently clear. This is because there are many different feature attribution methods, and different methods provide very different explanations (Bordt et al., 2022; Krishna et al., 2024). As such, it simply can’t be the case that all the different methods formalize the same intuitive concept. In the terminology introduced in Section 4.1, with feature attribution methods we have a situation not of convergence but of divergence.
Unfortunately, the lack of interpretation does not only apply to a few methods and not only to feature attributions. Take almost any recent paper that introduces a new explainability method or claims to improve an existing one. The abstract and introduction will usually start by noting how important explainability is. The paper itself will then be about some technical derivations and experiments. In the end, the paper claims to have contributed to explainability. However, the evidence for this is usually lacking. This is because the paper never discussed the matter of interpretation: Does the proposed method formalize an intuitive concept? Can the method contribute to answering any specific questions that a human decision-maker might have?
@@ 5.2 Statistics without interpretation lead to errors in interpretation
Many papers discuss failure modes of explanation algorithms (Adebayo et al., 2018; Kumar et al., 2020; Molnar et al., 2020; Bordt et al., 2022; Geirhos et al., 2023; Bove et al., 2024; Huang & Marques-Silva, 2024; Tomaszewska & Biecek, 2024). This literature is full of examples of invalid interpretations, for example
@@ 6 What Explainable Machine Learning can Learn from Applied Statistics
6.1 Instead of asking whether a model is “interpretable” or whether a given statistic is an “explanation”, we should ask what specific questions can be answered given a model or statistic
@@
Let us for a change illustrate this point with interpretable models, specifically the interpretable model class of Generalized Additive Models (GAMs) (Lou et al., 2013). GAMs represent the learned function as a set of univariate functions, where each function depicts the additive effect of one feature in the model. Given a GAM, certain specific questions can be directly answered by inspecting the model, whereas other questions cannot. Illustrative examples of questions that can be answered are “Are patients with pneumonia rated as high-risk?” and “How would the prediction change in response to a change in the input features?” (Caruana et al., 2015; Bosschieter et al., 2024). On the other hand, “Would the model discriminate against the members of a historically disadvantaged group in deployment?” is an example of a question that can not be answered by inspecting the interpretable model. In order to address this question, one would need to employ additional statistics, for example fairness metrics (Barocas et al., 2023). [^5] One of the advantages of interpretable model over post-hoc methods is that it is often fairly clear what specific questions can be answered given an interpretable model. See also Rudin (2019).
The key insight regarding future research in explainable machine learning is that papers should make explicit what specific questions a method is designed to answer (of course, this is directly related to our argument in Section 5).
6.2 Broad-based intuitive concepts like trust are unlikely to be formalized by explanations
In Section 4, we have discussed that research often starts with intuitive concepts and attempts to formalize them. This raises the question: What intuitive concepts can be successfully formalized? Of course, providing a general answer to this question is beyond the scope of our work. However, what stands out is that many of the intuitive concepts that motivate research in explainable machine learning are incredibly broad (consider, for example, the famous example of “trust” from the LIME paper (Ribeiro et al., 2016)). We would argue that a comparison with applied statistics clarifies that explainable machine learning is unlikely to succeed at formalizing concepts as broad as “trust” or “understanding” in individual statistics.
@@ 6.3 Explanations cannot replace other statistics such as fairness and robustness
Fairness and robustness measures are statistics of functions (Definition 2.3). This means they are fundamentally similar to explanations, even if often perceived differently. Moreover, fairness and robustness measures usually have an interpretation (Section 2).
What does our analysis imply for the relationship between these measures and explanations? Most importantly, it implies that it is not the purpose of explanations to determine fairness (even though this is sometimes argued in the literature, see, e.g. Slack et al. (2020)). This is because explanations should be designed to address specific questions (Section 6.1), and we already have statistics that formalize fairness notions. More generally, if we are interested in a particular property of a classifier and already have a statistic that formalizes it, then we should rely on this statistic instead of attempting to derive it implicitly through general-purpose explanations. This point is especially relevant for model auditing and regulation (Bordt et al., 2022; Cherian & Candès, 2024).
@@ 6.4 Working with explanations requires expertise
In the statistical sciences, it is well-recognized that working with statistics requires expertise (Tishkovskaya & Lancaster, 2012). For example, we educate students to understand experimental design and correctly interpret $p$-values and the $F$-statistic. Because explanations are at least as complex as these traditional statistics, we believe they will primarily be helpful to experts. Concretely, many currently popular methods, such as SHAP and gradient-based explanations, have complex technical properties and limitations related to technical notions such as in- and out-of distribution (Anders et al., 2020; Taufiq et al., 2023). Working with these explanations thus requires data science training. This does not necessarily have to be the case for all explanations — counterfactual explanations, in particular, can be interpreted without any specialized knowledge. Nevertheless, the analogy with applied statistics demonstrates that working with explanations, like other statistical quantities, requires expertise.
6.5 Benchmark datasets can help to probe explanations’ empirical properties but generally don’t provide interpretations
@@
At the same time, it is important to recognize that benchmarks are unlikely to provide interpretations for the output of post-hoc explanation algorithms. This is because interpretation is about the relationship between statistics and intuitive human concepts (see Section 4.1). To make this more concrete, let us again rely on the analogy between explainable machine learning and applied statistics and ask how interpretations of traditional statistics are obtained. At the risk of oversimplification, let us discuss the mean and the median. It is well-known that both the mean and the median are useful for interpreting probability distributions (for example, in the analysis of income distributions (Saez & Zucman, 2016)). At the same time, the limitations of these statistics are also clearly understood. For example, the mean is not robust to outliers, and neither statistic captures the skewness of the data. Now, the idea of collecting some representative probability distributions to benchmark the interpretative performance of the mean against the median subject to some externally defined metric seems quite absurd. The same is true for other statistics of probability distributions. By analogy, we argue that it is also true for statistics of functions.

### v5

# Position: Rethinking Explainable Machine Learning as Applied Statistics

@@ Abstract
In the rapidly growing literature on explanation algorithms, it often remains unclear what precisely these algorithms are for and how they should be used. In this position paper, we argue for a novel and pragmatic perspective: Explainable machine learning needs to recognize its parallels with applied statistics. Concretely, explanations are statistics of high-dimensional functions, and we should think about them analogously to traditional statistical quantities. Among others, this implies that we must think carefully about the matter of interpretation, or how the explanations relate to intuitive questions that humans have about the world. The fact that this is scarcely being discussed in research papers is one of the main drawbacks of the current literature. Moving forward, the analogy between explainable machine learning and applied statistics suggests fruitful ways for how research practices can be improved.
Explainable Machine Learning, Position Paper, ICML
@@ 1 Introduction
In this position paper, we argue for a novel perspective on explainable machine learning: The field needs to recognize its parallels with applied statistics. By applied statistics, we mean the use of statistical theory, methods, and tools to analyze data and solve real-world problems (Wilcox & Rousselet, 2023; Angrist & Pischke, 2009; Cox & Donnelly, 2011; Cox, 2018; Franconeri et al., 2021). Both explainable machine learning and applied statistics aim to summarize the behaviour of high-dimensional objects — functions in explainability, and datasets and their respective probability distributions in statistics (Wasserman, 2004). Moreover, both fields share a common goal: determining how relevant questions about the real world can be answered using such summaries.
@@
    The matter of its interpretation.
Mathematically, explanation algorithms have the form of a functional — a function of functions (Rudin, 1991). They take a function and map it to a simpler object, often a vector. One might say that the functional “explains” a function by reducing its complexity. However, not all dimension-reducing functionals have explanatory value. For this to be the case, the functional also needs to formalize an intuitive concept, a nuanced matter debated in the philosophy of science (Justus, 2012). If this is the case, then we say that the functional has an interpretation. Traditional statistical quantities like the $p$-value, confidence intervals and visualizations usually have an interpretation.
In the current literature on explainable machine learning, papers often introduce novel explanation algorithms without discussing their interpretation. This is problematic in itself because “explanations” without an interpretation miss the point. It is additionally problematic since historically, even relatively simple statistics that have an interpretation have frequently been misinterpreted, especially by applied scientists and decision-makers (Wasserstein & Lazar, 2016). Consequently, it is not surprising that the outputs of explanation algorithms have frequently been misinterpreted as well (Molnar et al., 2020).
@@
To be concise, this paper focuses on post-hoc explanation algorithms, which comprise the bulk of the research in explainable machine learning over the last decade. However, the analogy between explainable machine learning and applied statistics applies broadly. In Section 7, we briefly discuss how it applies to mechanistic interpretability.
@@ 2 Post-Hoc Explanation Algorithms are Statistics of Functions
Notation. $\mathcal{D}$ denotes a probability distribution on $\mathbb{R}^{d}$ and $f:\mathbb{R}^{d}\to\mathbb{R}$ a function. We let $\mathcal{P}(\mathbb{R}^{d})$ denote the space of all probability distributions on $\mathbb{R}^{d}$ with Borel $\sigma$-algebra and $\mathcal{F}(\mathbb{R}^{d})$ the space of all functions $f:\mathbb{R}^{d}\to\mathbb{R}$.
@@ 2.2 Explanation Algorithms: Statistics of functions
    SHAP values (Lundberg & Lee, 2017):
    $\scriptstyle\mathnormal{F}_{i}(f,\mathcal{D},x)=\sum_{S\subseteq[d]\setminus\{i\}}{\frac{|S|!(d-|S|-1)!}{|d|!}}[\mathbb{E}_{\mathcal{D}}(f|x_{S\cup\{i\}})-\mathbb{E}_{\mathcal{D}}(f|x_{S})].$
@@
    The closest point to $x$ with a different label than $x$:
    $\operatorname*{arg\,min}_{z\in\{y|f(y)\neq f(x)\}}||x-z||_{1}.$
@@
In summary, Definition 2.3 captures most, if not all, local post-hoc explanation algorithms. It also captures various other properties of functions, such as fairness- and robustness metrics. Note that when we talk about an algorithm, we mean the function implemented through that algorithm.
The fact that explanations are functionals of the learned predictor has previously been noted in the literature (Fisher et al., 2019; Marx et al., 2023; Scholbeck et al., 2024). In this work, we build on this insight to discuss the relationship between the mathematical aspects of explanation algorithms and their interpretation.
@@ 3 The Analogy Between Explainable Machine Learning and Applied Statistics
We have seen that post-hoc explanations and other properties of classifiers can be seen as statistics of functions and that there is a formal similarity between statistics of functions and statistics of probability distributions and datasets. This suggests that we should think about explanations and traditional statistical quantities along analogous lines and that explainable machine learning should be considered statistics for high-dimensional functions.
Historically, researchers in statistics have studied how the properties of probability distributions and large datasets can be summarized using distribution and dataset statistics (Stigler, 2002). In explainable machine learning, researchers have studied how the properties of learned functions can be summarized using explanations. In both cases, the principle is the same: We take a high-dimensional object, map it to a low-dimensional space, and then use the resulting statistical value(s) to answer relevant questions that we are interested in.
As we argue in the following Section 4, the analogy between explainable machine learning and statistics extends towards the crucial question of how to interpret explanations in applications. Before this, we want to briefly discuss two important insights.
@@
Insight 2: A large part of the literature in explainable machine learning studies the mathematical and computational properties of statistics of functions.
The mathematical framework introduced in Section 2 can help us to understand and categorize research questions in explainable machine learning. In particular, we can ask whether a research question can be fully expressed within the mathematical framework. In other words: Which research in explainable machine learning studies the mathematical properties of high-dimensional functionals? It turns out that a large part of the literature studies such questions. For example, many papers study the sensitivity of a statistic $F$ with respect to $f$ and $x$ (Slack et al., 2020; Anders et al., 2020; Lakkaraju et al., 2020), or the behaviour of a statistic $F$ on a subset of functions $\mathcal{I}\subset\mathcal{F}(\mathbb{R}^{d})$ (Garreau & von Luxburg, 2020; Bordt & von Luxburg, 2023). Similarly, many papers study the computational aspects of statistics of functions, for example, of Shapley values (Jethani et al., 2021; Covert et al., 2023; Wang et al., 2025).
Research on the mathematical properties of statistics can be useful. In particular, research in statistics also has strong mathematical aspects (Larsen & Marx, 2005). However, it is also important to highlight what this research usually does not discuss. Namely, what are the real-world questions that can be answered with the explanations, and why?
@@ 4 Statistics and their Interpretation
Both in applied statistics and explainable machine learning, the starting points are intuitive questions humans have about the real world or the model (Liao et al., 2020). For example,
@@ 4.1 Formalizing intuitive concepts
How can we determine whether a given statistic has an interpretation? It turns out that the relationship between intuitive concepts and their formalizations has long been studied in philosophy, where this kind of problem is referred to as conceptual analysis (Chalmers, 1996; Chalmers & Jackson, 2001; Jackson, 1998; Justus, 2012). A central insight in this literature is that the intuitive concept is informal and pre-theoretic, whereas the formalized concept is theoretic and part of a workable theory (Carnap, 1945, 1950). Since the intuitive concept is not formalized, there is no strict way that we could prove that a statistic formalizes an intuitive concept. However, there are arguments for whether a formalization captures an intuitive concept that come close to proving such statements.
@@ 4.2 Statistics of functions that have an interpretation
Examples of post-hoc explanation methods that have an interpretation are counterfactual explanations (Wachter et al., 2017) and anchors (Ribeiro et al., 2018). Counterfactual explanations formalize the intuitive concept “What would need to change to get a different prediction?”. Anchors formalize the intuitive concept that “Similar observations should be treated similarly”. What is important is that for both counterfactual explanations and anchors, it is clear which intuitive questions can and cannot be answered given the explanations. For example, a counterfactual explanation applies only to an individual data point and does not necessarily tell us how the input features of other points would need to be changed. Similarly, anchors only apply to a specific region of the input space.
In addition, the literature on the intersection of statistics and machine learning has introduced various notions of feature importance that clearly define what they mean by “important” (Strobl et al., 2008; Zhang & Janson, 2020; Verdinelli & Wasserman, 2024). This includes model reliance (Donnelly et al., 2023), conditional model reliance (Fisher et al., 2019), and a variety of other notions (Lei et al., 2018). For an introduction, see Ewald et al. (2024). While the correct interpretation of the respective feature importance scores can sometimes be quite technical (compare Section 6.4), the papers in this literature take the matter of interpretation seriously. As such, they present positive examples of statistics of functions that have an interpretation. Interestingly, these notions of feature importance usually concern the global importance of the variable (Covert et al., 2020).
@@ 4.3 Many post-hoc explanations lack an interpretation
As an example, consider local feature attribution methods as they are being discussed in the literature on explainable machine learning (Ribeiro et al., 2016; Lundberg & Lee, 2017; Selvaraju et al., 2017). These methods address the question, “Which features are important for the prediction?”. As such, there is an intuitive concept that the different methods attempt to formalize. However, the problem is that the relationship between formalism and intuitive concept is not sufficiently clear. This is because there are many different feature attribution methods, and different methods provide very different explanations (Bordt et al., 2022; Krishna et al., 2024). Consequently, it simply cannot be the case that all the different methods formalize the same intuitive concept. Applying the terminology introduced in Section 4.1, we observe that feature attribution methods do not exhibit convergence, but divergence.
Unfortunately, the lack of interpretation does not only apply to a few methods and not only to feature attributions. Take almost any recent paper in the machine learning literature that introduces a new explainability method or claims to improve an existing one. The abstract and introduction will usually start by noting how important explainability is. The paper itself will then be about some technical derivations and experiments. In the end, the paper claims to have contributed to explainability. However, the evidence for this is usually lacking. This is because the paper never discussed the matter of interpretation: Does the proposed method formalize an intuitive concept? Can the method contribute to answering any specific questions that a human decision-maker might have?
@@ 5.2 Statistics without interpretation lead to errors in interpretation
Many papers discuss failure modes of explanation algorithms (Adebayo et al., 2018; Kumar et al., 2020; Molnar et al., 2020; Bordt et al., 2022; Geirhos et al., 2024; Bove et al., 2024; Huang & Marques-Silva, 2024; Tomaszewska & Biecek, 2024). This literature is full of examples of invalid interpretations, for example
@@ 6 What Explainable Machine Learning can Learn from Applied Statistics
6.1 Instead of asking whether a model is “interpretable” or whether a given statistic is an “explanation”, we should ask which specific questions can be answered given a model or statistic
@@
Let us for a change illustrate this point with interpretable models, specifically the interpretable model class of Generalized Additive Models (GAMs) (Lou et al., 2013). GAMs represent the learned function as a set of univariate functions, where each function depicts the additive effect of one feature in the model. Given a GAM, certain specific questions can be directly answered by inspecting the model, whereas other questions cannot. Illustrative examples of questions that can be answered are “Are patients with pneumonia rated as high-risk?” and “How would the prediction change in response to a change in the input features?” (Caruana et al., 2015; Bosschieter et al., 2024). On the other hand, “Would the model discriminate against the members of a historically disadvantaged group in deployment?” is an example of a question that can not be answered by inspecting the interpretable model. In order to address this question, one would need to employ additional statistics, for example fairness metrics (Barocas et al., 2023). [^4] One of the advantages of interpretable model over post-hoc methods is that it is often fairly clear which specific questions can be answered given an interpretable model. See also Rudin (2019).
The key insight regarding future research in explainable machine learning is that papers should make explicit which specific questions a method is designed to answer (of course, this is directly related to our argument in Section 5).
6.2 Multifaceted intuitive concepts like trust are unlikely to be formalized by explanations
In Section 4, we have discussed that research often starts with intuitive concepts and attempts to formalize them. This raises the question: Which intuitive concepts can be successfully formalized? Of course, providing a general answer to this question is beyond the scope of our work. However, what stands out is that many of the intuitive concepts that motivate research in explainable machine learning are incredibly broad (consider, for example, the famous example of “trust” from the LIME paper (Ribeiro et al., 2016)). We would argue that a comparison with applied statistics clarifies that explainable machine learning is unlikely to succeed at formalizing concepts as broad as “trust” or “understanding” in individual statistics.
@@ 6.3 Explanations cannot replace other statistics such as fairness and robustness
Fairness and robustness measures are statistics of functions (Definition 2.3). This means that they are fundamentally similar to explanations, even if often perceived differently. Moreover, fairness and robustness measures usually have an interpretation (Section 2).
What does our analysis imply for the relationship between these measures and explanations? Most importantly, it implies that it is not the purpose of explanations to determine fairness (even though this is sometimes suggested in the literature, see, e.g. Slack et al. (2020)). This is because explanations should be designed to address specific questions (Section 6.1), and we already have statistics that formalize fairness notions. More generally, if we are interested in a particular property of a classifier and already have a statistic that formalizes it, then we should rely on this statistic instead of attempting to derive it implicitly through general-purpose explanations. This point is especially relevant for model auditing and regulation (Bordt et al., 2022; Cherian & Candès, 2024).
@@ 6.4 Working with explanations requires expertise
In the statistical sciences, it is well-recognized that working with statistics requires expertise (Tishkovskaya & Lancaster, 2012). For example, we educate students to understand experimental design and correctly interpret $p$-values and the $F$-statistic. Because explanations are at least as complex as these traditional statistics, we believe that they will primarily be helpful to experts. Concretely, many currently popular methods, such as SHAP and gradient-based explanations, have complex technical properties and limitations related to technical notions such as in- and out-of distribution (Anders et al., 2020; Taufiq et al., 2023). Working with these explanations thus requires data science training. This does not necessarily have to be the case for all explanations — counterfactual explanations, in particular, can be interpreted without any specialized knowledge. Nevertheless, the analogy with applied statistics demonstrates that working with explanations, like other statistical quantities, requires expertise.
6.5 Benchmark datasets can help to probe explanations’ empirical properties but generally do not provide interpretations
@@
At the same time, it is important to recognize that benchmarks are unlikely to provide interpretations for the output of post-hoc explanation algorithms. This is because interpretation is about the relationship between statistics and intuitive human concepts (see Section 4.1). To make this more concrete, let us again rely on the analogy between explainable machine learning and applied statistics and ask how interpretations of traditional statistics are obtained. At the risk of oversimplification, let us discuss the mean and the median. It is well-known that both the mean and the median are useful for interpreting probability distributions (for example, in the analysis of income distributions (Saez & Zucman, 2016)). At the same time, the limitations of these statistics are also clearly understood. For example, the mean is not robust to outliers, and neither statistic captures the skewness of the data. Now, the idea of collecting some representative probability distributions to benchmark the interpretative performance of the mean against the median subject to some externally defined metric seems quite absurd. The same is true for other statistics of probability distributions. By analogy, we argue this is also true for statistics of functions.
7 What about Mechanistic Interpretability?
To be concise, this paper has focused chiefly on post-hoc explanation algorithms. However, we believe that the analogy between explainable machine learning and applied statistics extends to the entire field. As another example, let us briefly outline how it applies to the increasingly popular field of mechanistic interpretability (Saphra & Wiegreffe, 2024; Sharkey et al., 2025).
Mechanistic interpretability aims to reverse-engineer the inner workings of neural networks (Elhage et al., 2021). It analyzes patterns in model weights and activations (Nanda et al., 2023) using a variety of methods, for example, probes (Gurnee et al., 2023), or by training sparse auto-encoders (Lieberum et al., 2024; Karvonen et al., 2024).
What is the relationship between mechanistic interpretability and the arguments outlined in this paper? We would argue that the relationship between applied statistics and mechanistic interpretability is even more apparent than for post-hoc explanations. This is because many papers in mechanistic interpretability literally apply traditional statistical methods to model weights and activations. Consider, for example, Nanda et al. (2023), where a key technique is to show that the Fourier decomposition of model weight matrices is sparse.
More generally, the weights and activations of a neural network fall under the concept of a statistic of a function (Definition 2.3). The same holds for derived statistics such as dimensionality reduction techniques and probes. However, note that there is an interesting difference between post-hoc explanation methods and mechanistic interpretability: Whereas post-hoc explanations offer a fixed statistic to be applied to all functions, researchers in mechanistic interpretability often aim to find a statistic that can provide insights about a given function. Nevertheless, once a statistic is chosen, the questions that should be asked about the statistic are the same as for post-hoc explanations (Section 4): Does the statistic formalize any intuitive concepts? What concrete questions can be answered with the statistic? In some cases, for example, when a probe relates to an interpretable outcome (Karvonen et al., 2024), mechanistic interpretability has found good answers to these questions.